Suppose at location x = 10, households consume 120 units of housing and the monthly commuting...

  1. Suppose at location x = 10, households consume 120 units of housing and the monthly commuting cost per kilometer is $60. The slope of the housing-price curve is − $[____].
Answer #1

Say the Budget for Housing is B

then Housing Price = B- 60*10 = B-600

Housing Price per unit = (B-600)/120

At X = 10, Housing Price per unit = (B-600)/120

Say of X=12, Housing Price = B- 60*12 = B-720

Housing Price per unit = (B-720)/120

At X = 12, Housing Price per unit = (B-720)/120

Hence slope = ((B-720)/120 - (B-600)/120)/(12-10)

=(B-720-B+600)/(120*2)

=-120/240

=-1/2

Hence Slope = -0.5